OpenStudy (anonymous):

Determine whether the graph of y = x2 − 4x + 1 has a maximum or minimum point, then find the maximum or minimum value. Minimum; (2, -3) Maximum; (2, -3) Minimum; (-3, 2) Maximum; (-3, 2)

OpenStudy (cwrw238):

the x^2 is positive so it will have a minimum value

OpenStudy (cwrw238):

y = x^2 − 4x + 1 = (x - 2)^2 -4 + 1 = (x-2)^2 - 3 - so minimum is at (2, -3)

OpenStudy (cwrw238):

2nd option

OpenStudy (anonymous):

Wouldn't it be the first one? because its min not max? @cwrw238

OpenStudy (cwrw238):

oh sorry - yes you are right